OpenStudy (purexyz):

Write an equation of a line in slope intercept form that is parallel to x+4y=6 passes through (-8,4)

OpenStudy (anonymous):

y = mx+b, is slope intercept form. m is the slope, since you are looking for parallel, you would have the same slope for the line that passes through (-8, 4).. to find that use this formula, y-y1= M(x-x1).. then isolate the y

OpenStudy (anonymous):

understand?

OpenStudy (purexyz):

yeah

OpenStudy (anonymous):

first you need to change x+4y=6 in to y = mx+b, then use y-y1= M(x-x1) and plug the point x1, y1 from (-8, 4)..use the same slope you find for x+4y=6 to solve for (-8, 4)

OpenStudy (anonymous):

let me know if your stuck

OpenStudy (purexyz):

so would it be 4y-4=(x-(-8))

OpenStudy (anonymous):

no when u changed x+4y=6 in to y = mx+b, it becomes